OpenStudy (anonymous):

Solve the system of equations and give the value x: 5x + 2y = 18 2x + 2y = 30

OpenStudy (anonymous):

do you know how to use the elimination method?

OpenStudy (anonymous):

No math is my weakest subject

OpenStudy (anonymous):

ok then 5x + 2y = 18 2x + 2y = 30 ------------- just change one of any equation into all negatives ex. 2x + 5y = 4 6x + 5y = 7 ----------- 2x + 5y = 4 -1(6x + 5y = 7) so can you try it by your self like how i did with the example?

OpenStudy (anonymous):

Doesn't make sense. Thanks though

OpenStudy (anonymous):

ok what dont you understand because i have to teach you not give you answers

OpenStudy (anonymous):

5x + 2y = 18 2x + 2y = 30 ------------- 5x + 2y = 18 -1(2x + 2y = 30) ------------- 5x + 2y = 18 -2x - 2y = -30 ------------- x = ?